xx GRHayLHDX

GRHayLHDX

Samuel Cupp Leonardo Rosa Werneck Terrence Pierre Jacques

Zachariah Etienne

Date

Abstract

Provides a hydrodynamic evolution code using GRHayLvia the GRHayLibthorn. This is a trimmed-down version of the IllinoisGRMHDthorn where magnetic fields are set to zero adapted to use the CarpetX driver thorn. The thorn currently only runs on cpus, though extension to gpus is planned.

1 Introduction

This thorn provides a GRHD evolution code built on the General Relativistic Hydrodynamics Library (GRHayL), which is included in the Einstein Toolkit via the GRHayLib thorn. The core library inherits or adapts code from the original IllinoisGRMHD thorn into a modular, infrastructure-agnostic library. While the IllinoisGRMHD thorn can be set to have zero magnetic fields, they will still be evolved. This is much slower than not evolving them, and GRHayLHD provides an implementation for evolving systems without magnetic fields.

This thorn currently only provides support for hybrid EOS, but work towards providing full support for tabulated EOS is ongoing. The core functions all come from GRHayLib, ensuring that any improvements or updates to the library can be easily adopted by the thorn. The conservative-to-primitive solver is selected at runtime, allowing for users to take advantage of any Con2Prim methods provided by

GRHayL.

2 Parameters

Most of the behavior of this thorn is controlled by GRHayLib, the thorn that provides GRHayL functionality within the Einstein Toolkit. However, there are several parameters in the thorn which control diagnostic or debugging features. Most are self-explanatory, but some parameters deserve a more detailed explanation.

First, many analysis or diagnostic thorns use HydroBase variables. By default, GRHayLHD never copies the data back from GRHayLHD variables to HydroBase variables, which would prevent the usage of these thorns. The ConvertxtoxHydroBasexevery parameter sets how often to copy back this data. Naturally, copying more frequently will slow down the simulation, so this should be set on the same frequency as the analysis or IO thorn that is using the HydroBase variable. We reduced the amount of data copying in GRHayLHD by using the HydroBase variables directly, which removes the need to copy them back and removes extra grid functions from memory. We still use distinct velocity variables because we use a different velocity than is defined in

HydroBase.

Next, there are several parameters for debugging or other testing. The perturbing parameters give users the ability to affect either the initial data or evolution data at runtime without slowing down standard simulations. The perturbxinitialxdata parameter controls whether the initial data from HydroBase is perturbed, and the perturbxeveryxcon2prim parameter controls whether the conservative variables are perturbed before every con2prim call. The magnitude of these perturbations is set by randomxpert.

3 Parameters

11 convertxtoxhydrobasexevery 10 Scope: private01  INT

11 Description: How often to convert GRHayLHDX primitive variables to HydroBase (Valencia formulation) primitive variables? Needed for some ET-based diagnostics. NOT needed for pure GRHayLHDX runs.

11  Range  10              01  Default: (none)
11 0:*  10 zero (disable) or positive (every N iterations)

11 matterxbc  10 Scope: private01  KEYWORD

11 Description: Chosen Matter boundary condition

11 Range 10 01 Default: outflow
11 copy  10 Copy BCs
11 outflow  10 Copy BCs with additional outflow BC for velocities
11 frozen  10 Frozen boundaries

11

11 perturbxeveryxcon2prim 10 Scope: private01  BOOLEAN

11 Description: Random perturbation applied every time con2prim runs

11                       10              01  Default: no

11

11 perturbxinitialxdata  10 Scope: private01  BOOLEAN

11 Description: Random perturbation to initial data

11                       10              01  Default: no

11

11 randomxpert  10 Scope: private01  REAL

11 Description: Magnitude of random perturbations

11 Range 10 01 Default: (none)
11 *:*  10 Anything goes.

11

11 randomxseed  10 Scope: private01  INT

11 Description: Random seed for random, generally roundoff-level perturbation on initial data. Seeds srand(), and rand() is used for the RNG.

11  Range  10              01  Default: (none)
11 0:99999999  10 Anything unsigned goes.

11

11 updatextmunu  10 Scope: private01  BOOLEAN

11 Description: Update Tmunu, for RHS of Einstein’s equations?

11                       10              01  Default: yes

11

11 verbose  10 Scope: private01  KEYWORD

11 Description: Determines how much evolution information is output

11 Range 10 01 Default: yes
11 no  10 Complete silence
11 yes  10 ”Outputs information about conservative-to-prim itive fixes, failures, etc.”

11

11 eosxtype  10 Scope: shared from GRHAYLIB01KEYWORD

11

11 evolvexentropy  10 Scope: shared from GRHAYLIB01BOOLEAN

4 Interfaces

General Implements: grhaylhdx

Inherits: admbasex tmunubasex hydrobasex grhaylib

Grid Variables

4.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  cccxspacetime  10                    01  centering 11 centering={1 1 1}
11                    10  cccxlapse  01  compact 11 0
11                    10  cccxbetax  01 description 11 Cell-centered spacetime quantities
11                    10  cccxbetay  01 dimensions 11 3
11                    10  cccxbetaz  01 distribution 11 DEFAULT
11                    10  cccxgxx  01 group type 11 GF
11                    10  cccxgxy  01  tags 11 InterpNumTimelevels=1 prolongation=”none” checkpoint=”no” parity={ +1 +1 +1 +1 +1 +1 -1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10  cccxgxz  01 timelevels 11 1
11                    10  cccxgyy  01 variable type 11 REAL
11  cccxcurv  10                    01  centering 11 centering={1 1 1}
11                    10  cccxkxx  01  compact 11 0
11                    10  cccxkxy  01 description 11 Cell-centered extrinsic curvature
11                    10  cccxkxz  01 dimensions 11 3
11                    10  cccxkyy  01 distribution 11 DEFAULT
11                    10  cccxkyz  01 group type 11 GF
11                    10  cccxkzz  01  tags 11 InterpNumTimelevels=1 prolongation=”none” checkpoint=”no” parity={ +1 +1 +1 +1 +1 +1 -1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  cccxtmunu  10                    01  centering 11 centering={1 1 1}
11                    10  cccxTtt  01  compact 11 0
11                    10  cccxTtx  01 description 11 Cell-centered stress-energy tensor
11                    10  cccxTty  01 dimensions 11 3
11                    10  cccxTtz  01 distribution 11 DEFAULT
11                    10  cccxTxx  01 group type 11 GF
11                    10  cccxTxy  01  tags 11 InterpNumTimelevels=1 prolongation=”none” checkpoint=”no” parity={ +1 +1 +1 +1 +1 +1 -1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10  cccxTxz  01 timelevels 11 1
11                    10  cccxTyy  01 variable type 11 REAL
11 grmhdxvelocities 10                    01  centering 11 centering={1 1 1}
11                    10  vx  01  compact 11 0
11                    10  vy  01 description 11 Components of primitive three velocity vx. Note that vx is defined in terms of 4-velocity as: vx = ux/uˆ0 . Note that this definition differs from the Valencia formalism.
11                    10  vz  01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 InterpNumTimelevels=1 prolongation=”none” checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  u0  10  u0  01  centering 11 centering={1 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 4-velocity component uˆ0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 InterpNumTimelevels=1 prolongation=”none” checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 grmhdxconservatives 10                    01  centering 11 centering={1 1 1}
11                    10  rhoxstar  01  compact 11 0
11                    10  tau  01 description 11 Evolved hydrodynamic variables
11                    10  Stildex  01 dimensions 11 3
11                    10  Stildey  01 distribution 11 DEFAULT
11                    10  Stildez  01 group type 11 GF
11                    10                    01  tags 11 rhs=”grmhdxconservativesxrhs” dependents=”GRHayLHDX::grmhdxconservativesxrhs GRHayLHDX::fluxxx GRHayLHDX::fluxxy GRHayLHDX::fluxxz”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  entxstar  10  entxstar  01  centering 11 centering={1 1 1}

11                       10                    01  compact 11 0

11                       10                    01 description 11 conservative entropy variable W S
11                       10                    01 dimensions 11 3
11                       10                    01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 rhs=”entxstarxrhs” dependents=”GRHayLHDX::entxstarxrhs GRHayLHDX::entxfluxxx GRHayLHDX::entxfluxxy GRHayLHDX::entxfluxxz”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL
11  yexstar  10  Yexstar  01  centering 11 centering={1 1 1}

11                       10                    01  compact 11 0

11                       10                    01 description 11 conservative electron fraction variable W rho Yxe
11                       10                    01 dimensions 11 3
11                       10                    01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 rhs=”Yexstarxrhs” dependents=”GRHayLHDX::Yexstarxrhs GRHayLHDX::Yexfluxxx GRHayLHDX::Yexfluxxy GRHayLHDX::Yexfluxxz”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL
11 grmhdxconservativesxrhs 10                    01  centering 11 centering={1 1 1}
11                       10  rhoxstarxrhs  01  compact 11 0
11                       10  tauxrhs  01 description 11 Right-hand side for grmhdxconservatives variables
11                       10  Stildexxrhs  01 dimensions 11 3
11                       10  Stildeyxrhs  01 distribution 11 DEFAULT
11                       10  Stildezxrhs  01 group type 11 GF
11                       10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL
11  entxstarxrhs  10  entxstarxrhs  01  centering 11 centering={1 1 1}

11                       10                    01  compact 11 0

11                       10                    01 description 11 RHS for partialxt \tilde{S} equation
11                       10                    01 dimensions 11 3
11                       10                    01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL
11  yexstarxrhs  10  Yexstarxrhs  01  centering 11 centering={1 1 1}

11                       10                    01  compact 11 0

11                       10                    01 description 11 RHS for partialxt \tilde{Y}xe equation
11                       10                    01 dimensions 11 3
11                       10                    01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL
11  fluxxx  10                    01  centering 11 centering={0 1 1}
11                       10  rhoxstarxfluxxx 01  compact 11 0
11                       10  tauxfluxxx  01 description 11 Fluxes in x direction
11                       10  Sxxfluxxx  01 dimensions 11 3
11                       10  Syxfluxxx  01 distribution 11 DEFAULT
11                       10  Szxfluxxx  01 group type 11 GF
11                       10                    01  tags 11 prolongation=”none” checkpoint=”no”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  fluxxy  10                    01  centering 11 centering={1 0 1}
11                    10  rhoxstarxfluxxy 01  compact 11 0
11                    10  tauxfluxxy  01 description 11 Fluxes in y direction
11                    10  Sxxfluxxy  01 dimensions 11 3
11                    10  Syxfluxxy  01 distribution 11 DEFAULT
11                    10  Szxfluxxy  01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  fluxxz  10                    01  centering 11 centering={1 1 0}
11                    10  rhoxstarxfluxxz 01  compact 11 0
11                    10  tauxfluxxz  01 description 11 Fluxes in z direction
11                    10  Sxxfluxxz  01 dimensions 11 3
11                    10  Syxfluxxz  01 distribution 11 DEFAULT
11                    10  Szxfluxxz  01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  entxfluxxx  10  entxfluxxx  01  centering 11 centering={0 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Entropy flux in x direction
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  entxfluxxy  10  entxfluxxy  01  centering 11 centering={1 0 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Entropy flux in y direction
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  entxfluxxz  10  entxfluxxz  01  centering 11 centering={1 1 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Entropy flux in z direction
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  yexfluxxx  10  Yexfluxxx  01  centering 11 centering={0 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Electron fraction flux in x direction
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  yexfluxxy  10  Yexfluxxy  01  centering 11 centering={1 0 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Electron fraction flux in y direction
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  yexfluxxz  10  Yexfluxxz  01  centering 11 centering={1 1 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Electron fraction flux in z direction
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 failurexchecker 10  failurexchecker 01  centering 11 centering={1 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Gridfunction to track conservative-to-primitives solver fixes. Beware that this gridfunction is overwritten at each RK substep.
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” Checkpoint=”no” InterpNumTimelevels=1
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

Uses header: GRHayLib.h

5 Schedule This section lists all the variables which are assigned storage by thorn GRHayLET/GRHayLHDX. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 Conditional:
11 grmhdxconservatives grmhdxvelocities 11 Yexstar Yexstarxrhs Yexfluxxx Yexfluxxy Yexfluxxz
11 u0 grmhdxconservativesxrhs  11 entxstar entxstarxrhs entxfluxxx entxfluxxy entxfluxxz
11 fluxxx fluxxy fluxxz failurexchecker 11 x

11 x  11 x

Scheduled Functions

HydroBaseXxPostInitial (conditional) x grhaylhdxxprim2con2prim x initialize grhaylhdx variables using hydrobasex initial data x

11 x 11 Type: 11 group

11

GRHayLHDXxPrim2Con2Prim (conditional) x grhaylhdxxcomputexcccxcenteredxspacetimexquantities x interpolate spacetime quantities to cell centers x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbasex::metric
11 x 11 x  11 admbasex::lapse
11 x 11 x  11 admbasex::shift
11 x 11 Type:  11 function
11 x 11 Writes: 11 cccxspacetime(everywhere)

11

GRHayLHDXxCon2Prim (conditional) x grhaylhdxxconservsxtoxprims x compute primitive variables from conservatives x

11 x 11 After: 11 grhaylhdxxcomputexcccxcenteredxspacetimexquantities
11 x 11 Type: 11 group

11

TmunuBaseXxAddToTmunu (conditional) x grhaylhdxxcomputextmunu x compute stress-energy tensor and interpolate to vertices x

11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 grmhdxvelocities

11 x 11 x  11 u0

11 x 11 x  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::eps
11 x 11 Type:  11 function
11 x 11 Writes: 11 cccxtmunu(everywhere)
11 x 11 x  11 tmunubasex::ettt
11 x 11 x  11 tmunubasex::etti
11 x 11 x  11 tmunubasex::etij

11

ODESolversxRHS (conditional) x grhaylhdxxrhs x evaluate rhss grhd equations x

11 x 11 Type: 11 group

GRHayLHDXxRHS (conditional) x grhaylhdxxcomputexcccxcenteredxextrinsicxcurvature x interpolate extrinsic curvature to cell centers x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbasex::curv
11 x 11 Type:  11 function
11 x 11 Writes: 11 cccxcurv(everywhere)

11

GRHayLHDXxRHS (conditional) x grhaylhdxxevaluatexsourcesxrhs x evaluate source terms in grhd rhss x

11 x 11 After: 11 grhaylhdxxcomputexcccxcenteredxextrinsicxcurvature
11 x 11 Type: 11 group

11

GRHayLHDXxRHS (conditional) x grhaylhdxxevaluatexfluxes x evaluate fluxes for grhd rhss x

11 x 11 After: 11 grhaylhdxxevaluatexsourcesxrhs
11 x 11 Type: 11 group

11

GRHayLHDXxRHS (conditional) x grhaylhdxxevaluatexfluxesxrhs x evaluate flux terms in grhd rhss x

11 x 11 After: 11 grhaylhdxxevaluatexfluxes
11 x 11 Type: 11 group

11

GRHayLHDXxRHS (conditional) x grhaylhdxxvalidatexyestar x x

11 x 11 After:  11 grhaylhdxxevaluatexfluxesxrhs
11 x 11 Language: 11 c
11 x 11 Type:  11 function
11 x 11 Writes: 11 yexstar(everywhere)
11 x 11 x  11 yexstarxrhs(everywhere)

11

GRHayLHDXxRHS (conditional) x grhaylhdxxvalidatexentstar

11 x 11 After:  11 grhaylhdxxevaluatexfluxesxrhs
11 x 11 Language: 11 c
11 x 11 Type:  11 function
11 x 11 Writes: 11 entxstar(everywhere)
11 x 11 x  11 entxstarxrhs(everywhere)

11

GRHayLHDXxRHS (conditional) x convertxgrhaylhdxxtoxhydrobase x convert grhaylhdx variables to hydrobasex x

11 x 11 After:  11 grhaylhdxxevaluatexfluxesxrhs
11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobasex::vel(everywhere)

11

GRHayLHDXxPrim2Con2Prim (conditional) x convertxhydrobasextoxgrhaylhdx x convert hydrobasex initial data to grhaylhdx variables x

11 x 11 After:  11 grhaylhdxxcomputexcccxcenteredxspacetimexquantities
11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::vel
11 x 11 x  11 cccxspacetime
11 x 11 Sync:  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxvelocities(everywhere)

11

CCTKxANALYSIS (conditional) x grhaylhdxxcomputexcccxcenteredxspacetimexquantities x interpolate spacetime quantities to cell centers x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbasex::metric
11 x 11 x  11 admbasex::lapse
11 x 11 x  11 admbasex::shift
11 x 11 Type:  11 function
11 x 11 Writes: 11 cccxspacetime(everywhere)

11

CCTKxANALYSIS (conditional) x convertxgrhaylhdxxtoxhydrobase x convert grhaylhdx variables to hydrobasex

11 x 11 After:  11 grhaylhdxxcomputexcccxcenteredxspacetimexquantities
11 x 11 x  11 mlxbssnxevolcalcgroup
11 x 11 Before: 11 particlextraceret
11 x 11 x  11 volumeintegralgroup
11 x 11 x  11 convertxtoxmhdxx3velocity
11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobasex::vel(everywhere)

11

GRHayLHDXxprimsxtoxconservs (conditional) x grhaylhdxxhybridxentropyxprimsxtoxconservs x entropy+hybrid version of grhaylhdxxprimsxtoxconservs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::eps
11 x 11 x  11 hydrobasex::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)
11 x 11 x  11 hydrobasex::entropy(everywhere)

11

GRHayLHDXxconservsxtoxprims (conditional) x grhaylhdxxhybridxentropyxconservsxtoxprims x entropy+hybrid version of grhaylhdxxconservsxtoxprims x

11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)
11 x 11 x  11 hydrobasex::entropy(everywhere)

11

GRHayLHDXxevaluatexsourcesxrhs (conditional) x grhaylhdxxhybridxentropyxevaluatexsourcesxrhs x entropy+hybrid version of grhaylhdxxevaluatexsourcesxrhs

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::entropy
11 x 11 x  11 cccxspacetime
11 x 11 x  11 cccxcurv
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxrhs

11

GRHayLHDXxevaluatexfluxes (conditional) x grhaylhdxxhybridxentropyxevaluatexfluxes x entropy+hybrid version of grhaylhdxxevaluatexfluxes x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::entropy
11 x 11 x  11 cccxspacetime
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 fluxxx
11 x 11 x  11 fluxxy
11 x 11 x  11 fluxxz
11 x 11 x  11 entxfluxxx
11 x 11 x  11 entxfluxxy
11 x 11 x  11 entxfluxxz

11

GRHayLHDXxevaluatexfluxesxrhs (conditional) x grhaylhdxxhybridxentropyxevaluatexfluxesxrhs x entropy+hybrid version of grhaylhdxxevaluatexfluxesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 fluxxx(interior)
11 x 11 x  11 fluxxy(interior)
11 x 11 x  11 fluxxz(interior)
11 x 11 x  11 entxfluxxx(interior)
11 x 11 x  11 entxfluxxy(interior)
11 x 11 x  11 entxfluxxz(interior)
11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 x  11 entxstarxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxrhs

11

GRHayLHDXxperturbxprimitives (conditional) x grhaylhdxxhybridxentropyxperturbxprimitives x entropy+hybrid version of grhaylhdxxperturbxprimitives

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::eps
11 x 11 x  11 hydrobasex::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)
11 x 11 x  11 hydrobasex::entropy(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)

11

GRHayLHDXxperturbxconservatives (conditional) x grhaylhdxxhybridxentropyxperturbxconservatives x entropy+hybrid version of grhaylhdxxperturbxconservatives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)

11

GRHayLHDXxprimsxtoxconservs (conditional) x grhaylhdxxhybridxprimsxtoxconservs x hybrid version of grhaylhdxxprimsxtoxconservs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::eps
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)

11

GRHayLHDXxPrim2Con2Prim (conditional) x grhaylhdxxperturbxprimitives x perturb initial primitive data x

11 x 11 After: 11 convertxhydrobasextoxgrhaylhdx
11 x 11 Before: 11 grhaylhdxxprimsxtoxconservs
11 x 11 Type: 11 group

11

GRHayLHDXxconservsxtoxprims (conditional) x grhaylhdxxhybridxconservsxtoxprims x hybrid version of grhaylhdxxconservsxtoxprims x

11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 grmhdxconservatives
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)

11

GRHayLHDXxevaluatexsourcesxrhs (conditional) x grhaylhdxxhybridxevaluatexsourcesxrhs x hybrid version of grhaylhdxxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 cccxspacetime
11 x 11 x  11 cccxcurv
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs

11

GRHayLHDXxevaluatexfluxes (conditional) x grhaylhdxxhybridxevaluatexfluxes x hybrid version of grhaylhdxxevaluatexfluxes x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 cccxspacetime
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 fluxxx
11 x 11 x  11 fluxxy
11 x 11 x  11 fluxxz

11

GRHayLHDXxevaluatexfluxesxrhs (conditional) x grhaylhdxxhybridxevaluatexfluxesxrhs x hybrid version of grhaylhdxxevaluatexfluxesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 fluxxx(interior)
11 x 11 x  11 fluxxy(interior)
11 x 11 x  11 fluxxz(interior)
11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs

GRHayLHDXxperturbxprimitives (conditional) x grhaylhdxxhybridxperturbxprimitives x hybrid version of grhaylhdxxperturbxprimitives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::eps
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)

11

GRHayLHDXxperturbxconservatives (conditional) x grhaylhdxxhybridxperturbxconservatives x hybrid version of grhaylhdxxperturbxconservatives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)

11

GRHayLHDXxprimsxtoxconservs (conditional) x grhaylhdxxtabulatedxentropyxprimsxtoxconservs x entropy+tabulated version of grhaylhdxxprimsxtoxconservs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::eps
11 x 11 x  11 hydrobasex::entropy
11 x 11 x  11 hydrobasex::ye
11 x 11 x  11 hydrobasex::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)
11 x 11 x  11 hydrobasex::entropy(everywhere)
11 x 11 x  11 hydrobasex::ye(everywhere)
11 x 11 x  11 hydrobasex::temperature(everywhere)

11

GRHayLHDXxconservsxtoxprims (conditional) x grhaylhdxxtabulatedxentropyxconservsxtoxprims x entropy+tabulated version of grhaylhdxxconservsxtoxprims

11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)
11 x 11 x  11 hydrobasex::entropy(everywhere)
11 x 11 x  11 hydrobasex::ye(everywhere)
11 x 11 x  11 hydrobasex::temperature(everywhere)

11

GRHayLHDXxevaluatexsourcesxrhs (conditional) x grhaylhdxxtabulatedxentropyxevaluatexsourcesxrhs x entropy+tabulated version of grhaylhdxxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::entropy
11 x 11 x  11 hydrobasex::ye
11 x 11 x  11 hydrobasex::temperature
11 x 11 x  11 cccxspacetime
11 x 11 x  11 cccxcurv
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxrhs
11 x 11 x  11 yexstarxrhs

11

GRHayLHDXxevaluatexfluxes (conditional) x grhaylhdxxtabulatedxentropyxevaluatexfluxes x entropy+tabulated version of grhaylhdxxevaluatexfluxes x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::entropy
11 x 11 x  11 hydrobasex::ye
11 x 11 x  11 hydrobasex::temperature
11 x 11 x  11 cccxspacetime
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 fluxxx
11 x 11 x  11 fluxxy
11 x 11 x  11 fluxxz
11 x 11 x  11 entxfluxxx
11 x 11 x  11 entxfluxxy
11 x 11 x  11 entxfluxxz
11 x 11 x  11 yexfluxxx
11 x 11 x  11 yexfluxxy
11 x 11 x  11 yexfluxxz

GRHayLHDXxPrim2Con2Prim (conditional) x grhaylhdxxprimsxtoxconservs x enforce simulation limits on primitives and compute conservatives x

11 x 11 After: 11 convertxhydrobasextoxgrhaylhdx
11 x 11 Type: 11 group

11

GRHayLHDXxevaluatexfluxesxrhs (conditional) x grhaylhdxxtabulatedxentropyxevaluatexfluxesxrhs x entropy+tabulated version of grhaylhdxxevaluatexfluxesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 fluxxx(interior)
11 x 11 x  11 fluxxy(interior)
11 x 11 x  11 fluxxz(interior)
11 x 11 x  11 entxfluxxx(interior)
11 x 11 x  11 entxfluxxy(interior)
11 x 11 x  11 entxfluxxz(interior)
11 x 11 x  11 yexfluxxx(interior)
11 x 11 x  11 yexfluxxy(interior)
11 x 11 x  11 yexfluxxz(interior)
11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 x  11 entxstarxrhs(interior)
11 x 11 x  11 yexstarxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxrhs
11 x 11 x  11 yexstarxrhs

11

GRHayLHDXxperturbxprimitives (conditional) x grhaylhdxxtabulatedxentropyxperturbxprimitives x entropy+tabulated version of grhaylhdxxperturbxprimitives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::eps
11 x 11 x  11 hydrobasex::entropy
11 x 11 x  11 hydrobasex::ye
11 x 11 x  11 hydrobasex::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)
11 x 11 x  11 hydrobasex::entropy(everywhere)
11 x 11 x  11 hydrobasex::ye(everywhere)
11 x 11 x  11 hydrobasex::temperature(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)

11

GRHayLHDXxperturbxconservatives (conditional) x grhaylhdxxtabulatedxentropyxperturbxconservatives x entropy+tabulated version of grhaylhdxxperturbxconservatives

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)

11

GRHayLHDXxprimsxtoxconservs (conditional) x grhaylhdxxtabulatedxprimsxtoxconservs x tabulated version of grhaylhdxxprimsxtoxconservs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::eps
11 x 11 x  11 hydrobasex::ye
11 x 11 x  11 hydrobasex::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)
11 x 11 x  11 hydrobasex::ye(everywhere)
11 x 11 x  11 hydrobasex::temperature(everywhere)

11

GRHayLHDXxconservsxtoxprims (conditional) x grhaylhdxxtabulatedxconservsxtoxprims x tabulated version of grhaylhdxxconservsxtoxprims x

11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)
11 x 11 x  11 hydrobasex::ye(everywhere)
11 x 11 x  11 hydrobasex::temperature(everywhere)

11

GRHayLHDXxevaluatexsourcesxrhs (conditional) x grhaylhdxxtabulatedxevaluatexsourcesxrhs x tabulated version of grhaylhdxxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::ye
11 x 11 x  11 hydrobasex::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 cccxspacetime
11 x 11 x  11 cccxcurv
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 yexstarxrhs

11

GRHayLHDXxevaluatexfluxes (conditional) x grhaylhdxxtabulatedxevaluatexfluxes x tabulated version of grhaylhdxxevaluatexfluxes x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::ye
11 x 11 x  11 hydrobasex::temperature
11 x 11 x  11 cccxspacetime
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 fluxxx
11 x 11 x  11 fluxxy
11 x 11 x  11 fluxxz
11 x 11 x  11 yexfluxxx
11 x 11 x  11 yexfluxxy
11 x 11 x  11 yexfluxxz

11

GRHayLHDXxevaluatexfluxesxrhs (conditional) x grhaylhdxxtabulatedxevaluatexfluxesxrhs x tabulated version of grhaylhdxxevaluatexfluxesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 fluxxx(interior)
11 x 11 x  11 fluxxy(interior)
11 x 11 x  11 fluxxz(interior)
11 x 11 x  11 yexfluxxx(interior)
11 x 11 x  11 yexfluxxy(interior)
11 x 11 x  11 yexfluxxz(interior)
11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 x  11 yexstarxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 yexstarxrhs

11

GRHayLHDXxperturbxprimitives (conditional) x grhaylhdxxtabulatedxperturbxprimitives x tabulated version of grhaylhdxxperturbxprimitives

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobasex::rho
11 x 11 x  11 hydrobasex::press
11 x 11 x  11 hydrobasex::eps
11 x 11 x  11 hydrobasex::ye
11 x 11 x  11 hydrobasex::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobasex::rho(everywhere)
11 x 11 x  11 hydrobasex::press(everywhere)
11 x 11 x  11 hydrobasex::eps(everywhere)
11 x 11 x  11 hydrobasex::ye(everywhere)
11 x 11 x  11 hydrobasex::temperature(everywhere)

11

GRHayLHDXxperturbxconservatives (conditional) x grhaylhdxxtabulatedxperturbxconservatives x tabulated version of grhaylhdxxperturbxconservatives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 yexstar(everywhere)

11

GRHayLHDXxPrim2Con2Prim (conditional) x grhaylhdxxconservsxtoxprims x compute primitive variables from conservatives x

11 x 11 After: 11 grhaylhdxxprimsxtoxconservs
11 x 11 Type: 11 group

11

GRHayLHDXxPrim2Con2Prim (conditional) x convertxgrhaylhdxxtoxhydrobase x convert grhaylhdx variables to hydrobasex x

11 x 11 After:  11 grhaylhdxxconservsxtoxprims
11 x 11 Language: 11 c
11 x 11 Reads:  11 cccxspacetime
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobasex::vel(everywhere)

11

ODESolversxPostStep (conditional) x grhaylhdxxcon2prim x compute primitive variables from conservatives x

11 x 11 Type: 11 group

GRHayLHDXxCon2Prim (conditional) x grhaylhdxxcomputexcccxcenteredxspacetimexquantities x interpolate spacetime quantities to cell centers x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbasex::metric
11 x 11 x  11 admbasex::lapse
11 x 11 x  11 admbasex::shift
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 x  11 entxstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 cccxspacetime(everywhere)

11

GRHayLHDXxCon2Prim (conditional) x grhaylhdxxperturbxconservatives x perturb conservatives before con2prim x

11 x 11 After: 11 grhaylhdxxcomputexcccxcenteredxspacetimexquantities
11 x 11 Before: 11 grhaylhdxxconservsxtoxprims
11 x 11 Type: 11 group